OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3.

OpenStudy (dtan5457):

Are any of those numbers powers, or just regular numbers?

OpenStudy (anonymous):

I can do it

OpenStudy (anonymous):

the ones after the variables are powers

OpenStudy (anonymous):

-3(2)2 - (-1)3 + 3(3)2 - 2(-1)3 -6(2) + 3 + 9(2) +2(3) -12 + 3 + 18+ 6

OpenStudy (anonymous):

-9 + 24

OpenStudy (anonymous):

Now i cant give direct answers so whats your answer

OpenStudy (anonymous):

is it 15

OpenStudy (dtan5457):

\[-3(2)^2-(-1)^3+3(3)^2-2(-1)^3\]

OpenStudy (anonymous):

yep

OpenStudy (dtan5457):

pemdas and all that

OpenStudy (anonymous):

Medal and fan :)

OpenStudy (anonymous):

thanks boo

OpenStudy (dtan5457):

@XxH2O_AvengerxX it said the numbers after the variables are powers

OpenStudy (anonymous):

ohhhh so there exponents??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh...lemme do it over real quick ok ? xD

OpenStudy (dtan5457):

ok

OpenStudy (anonymous):

-6^2 - (-1)^3 +3(3)^2 + 2(-1)^3

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

-12 + 3 + 9^2 + -2^3 -12 +3 + 81 - 8 9 + 73

OpenStudy (anonymous):

is it 82

OpenStudy (dtan5457):

@XxH2O_AvengerxX you completely forgot about PEMDAS lol

OpenStudy (dtan5457):

exponents first, then multiplication wanna try one more time?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

-.- ugh fine

OpenStudy (anonymous):

dtan why dont you try it then

OpenStudy (anonymous):

-6^2 - (-1)^3 +3(3)^2 + 2(-1)^3 I GOT THIS

OpenStudy (dtan5457):

He started it first, I want him to finish it lol unless he can't do it, which I'm sure he acn

OpenStudy (dtan5457):

*can

OpenStudy (anonymous):

wait @dtan5457 , so does -6^2 turn to 36? Doesnt it turn to a postitve?

OpenStudy (dtan5457):

\[-3(2)^2=-3(4)\]

OpenStudy (dtan5457):

like i said, exponents first

OpenStudy (dtan5457):

-6^2 is not part of the question but that would be -36. (-6)^2=36

OpenStudy (dtan5457):

parenthesis make a difference

OpenStudy (anonymous):

-6^2 - (-1)^3 +3(3)^2 + 2(-1)^3 36 + 3 + 3(9) + 2(3) 36 + 3 + 27 + 6 39 + 33

OpenStudy (anonymous):

Is that it??? :(

OpenStudy (dtan5457):

no lol

OpenStudy (anonymous):

Ugh!!!!! I'm done then

OpenStudy (anonymous):

oh my goodness can someone help me please

OpenStudy (anonymous):

\(\huge\frak\color{#FF1FF1}{Im~gonna~stick~to~L^a_T~^e_X}\)

OpenStudy (dtan5457):

\[-3(2)^2-(-1)^3+3(3)^2-2(-1)^3=-3(4)-(-1)+27+2\]

OpenStudy (dtan5457):

@XxH2O_AvengerxX Do you see the mistake?

OpenStudy (anonymous):

Yeah now i do.....

OpenStudy (dtan5457):

yeah, well i'm gonna medal you for effort regardless lol

OpenStudy (dtan5457):

@tempboo I simplified it for you, now just do the basic operations

OpenStudy (anonymous):

im still confused

OpenStudy (dtan5457):

\[−3(4)−(−1)+27+2\] Solve this, all exponents are gone now, should be easy

OpenStudy (anonymous):

its just a whole bunch of question marks u sent

OpenStudy (dtan5457):

refresh

OpenStudy (anonymous):

is it 18

OpenStudy (anonymous):

Yep..........